# Infusion: internal diffusion for inpainting of dynamic textures and complex motion

### v3

@@ 3.3. Single Video (Internal) Training
3.4. Interval training
@@
To circumvent this problem, we propose to use one lightweight network, trained on one subset of timesteps at a time. The training starts from the end of the diffusion process on the first interval. Once training is done on this interval, we use the model to perform the inference on the test data for all timesteps of the interval, up to the beginning of the next interval. The model is then trained on the next time interval. This is carried out until we have reached time step $t=0$. We can, of course, save the intermediate models if we wish, but this is not really necessary for video inpainting, since the model can only be used for the current video.
@@
Table 1. Inpainting metrics on dynamic texture datasets. Our methods perform the best in terms of LPIPS and SVFID. Interval training brings a noticeable improvement over the baseline approach to diffusion. Among the recent methods, only our method is able to produce diverse results.
@@
LPIPS ↓ SVFID ↓ PSNR (dB) ↑ SSIM ↑ Diversity ↑ LPIPS ↓ SVFID ↓ PSNR (dB) ↑ SSIM ↑
Patch (Newson et al., 2014) 0.051 0.227 29.47 0.954 0.053 0.0180 0.055 36.52 0.988
ISVI (Zhang et al., 2022c) 0.062 0.500 27.41 0.945 0 0.0302 0.248 34.20 0.982
E2FGVI (Li et al., 2022) 0.046 0.288 31.58 0.961 0 0.0145 0.078 38.76 0.991
FGT (Zhang et al., 2022b) 0.046 0.202 30.88 0.956 0 0.0144 0.059 37.99 0.990
ProPainter (Zhou et al., 2023a) 0.046 0.261 31.77 0.960 0 0.0137 0.051 39.62 0.991
Infusion (no interval) 0.042 0.145 29.54 0.960 0.039 0.0150 0.043 37.09 0.989
@@
This strategy has several advantages. First, we can use a lighter network as it specializes only on a small subset of timesteps. Second, we can avoid complicated weighting schemes for the loss function: in one interval, all timesteps have roughly the same weight and the weighting term can thus be set to $1$. This is not the case when one network is used and must perform well on all $T$ timesteps (Hang et al., 2023; Nichol and Dhariwal, 2021).
4. Experiments
4.1. Evaluation and masks
We compare our method with several state of the art methods. We include a classical patch-based approach (Newson et al., 2014) which is also internal and well adapted to the inpainting of dynamic textures, as well as more recent deep learning approaches: E2FGVI (Li et al., 2022), ISVI (Zhang et al., 2022c), FGT (Zhang et al., 2022b), and ProPainter (Zhou et al., 2023a).
We evaluate our method on complex videos featuring dynamic textures, large occlusions, and moving objects. These include, firstly, video sequences with a segmented object to be removed, from (Newson et al., 2014; Granados et al., 2012; Le et al., 2019), as well as some new sequences also with segmented objects. Secondly, we use dynamic texture databases: (Tesfaldet et al., 2018), featuring 59 textures, and a subset of DTDB (Hadji and Wildes, 2018) containing one random video from each appearance category. In these last two databases, we use synthetic masks, since there are no meaningful objects (people walking, etc) to be removed.
@@ 4.1. Evaluation and masks
A database which is often used to evaluate generic video inpainting methods, not included in our evaluation, is the DAVIS database (Perazzi et al., 2016). This database was initially developed to benchmark video segmentation methods. It contains a large amount of videos with strong camera motion, for which the optical flow is of crucial importance. On this database, methods combining optical flow and deep learning are clearly superior to ours, which is not adapted to such situations.
@@
Figure 4. In these scenes, our method satisfactorily inpaints dynamic textures preserving sharp details (top). It also better handles complex occlusions such as people crossing paths.
4.2. Training details
@@
For training, we generate masks using random strokes similarly to Kim et al. (2019). Training masks may overlap with the test masks, in which case no loss is actually computed in the intersection.
We use the Adam optimizer with an initial learning rate of 1e-4, and train for 200k total steps. For interval training, the 1000 timesteps are divided into 20 intervals of length 50. The same training budget is allocated to each interval.
4.3. Quantitative results
We quantitatively evaluate the methods using Peak Signal-to-Noise Ratio (PSNR), structure similarity measure (SSIM), Learned Perceptual Image Patch Similarity (Zhang et al., 2018) (LPIPS) and Single Video Fréchet Inception Distance (SVFID) (Wang et al., 2018; Gur et al., 2020). Finally, we also compute a diversity score following (Shaham et al., 2019). As explained above, the evaluation is conducted on the two databases of dynamic textures.
The first three metrics (PSNR, SSIM and LPIPS) are computed frame-by-frame and therefore do not reflect motion discontinuities. SVFID, on the other hand, takes spatio-temporal features into account. For computing the SVFID, we follow (Dorkenwald et al., 2021) and use the same model, trained on dynamic textures, for extracting the spatio-temporal features required for its computation.
Results are given in Table 1. We first observe that the LPIPS metric is significantly better with our method. In contrast, PSNR and SSIM are relatively good for all methods, the SSIM being slightly better for us and the PSNR better for the two deep-learning based methods (Zhou et al., 2023a)and (Li et al., 2022). A visual inspection of the videos provided with this paper nevertheless clearly shows that such methods struggle to reproduce spatial details of the textures. This can be explained by the fact that the PSNR is well-known to be poorly adapted to fine details (Blau and Michaeli, 2018; Saharia et al., 2022b). The LPIPS metric, on the other hand, is more faithful to human perception.
@@ 4.3. Quantitative results
The diversity score shows a superiority of our approach and of the patch-based method from (Newson et al., 2014) with respect to other approaches. In fact, all considered deep-learning methods always produce the same result, yielding a null diversity score.
@@
Table 2. Inpainting metrics when averaging 100 samples or using a single one on the Tesfaldet dataset. PSNR is improved by averaging but the perceptual metrics are severely degraded.
[2] Method LPIPS ↓ SVFID ↓ PSNR (dB) ↑ SSIM ↑ Diversity ↑ Infusion (avg) 0.056 0.231 30.97 0.962 0 Infusion 0.035 0.116 30.27 0.964 0.054
4.4. Object removal experiments
@@
In this section, we consider the problem of object removal through several classical sequences used in previous works on video inpainting (Newson et al., 2014; Granados et al., 2012; Le et al., 2019) and a few new ones. We consider the same methods as before. For each sequence, a mask is given corresponding to an object (often a person) to be removed. As already mentioned, there is no ground truth for these sequences, so that quantitative evaluation is not performed. Nevertheless, such situations correspond to real applications and a more realistic foreground-background articulation than with synthetic masks. All results can be found in the supplementary materials. One can observe a much better respect of backgrounds such as dynamic textures or occluded moving objects with our approach than with other deep-learning based ones. This is briefly illustrated in Figure 4. On the first line, one sees that the unstable reflection of the building in the water is better reproduced by our methods than by FGT and ProPainter. The second and third rows display cases with occluded moving objects. In this case, results are blurry for ProPainter and blending artifacts are visible for FGT, while ours yields more satisfactory structures. The reader is strongly advised to look at the videos in the supplementary materials. These videos support the claim that a diffusion framework with appropriate internal learning is a good choice for stochastic and complex content.
4.5. Ablation study
@@
This section demonstrates the importance of interval training. We compare the baseline training (i.e. training on all timesteps) and interval training. We train a neural network with the same number of parameters and for the same total number of training steps for both approaches. Two visual results can be seen in Figure 5. We observe that interval training clearly improves the reproduction of fine, textured details. More results can be seen on the videos in the supplementary materials. This observation is confirmed quantitatively in Table 1. Interval training significantly improves the quantitative results, especially for the perceptual metrics (LPIPS, SVFID) as well as for diversity.
@@
We emphasize that in the baseline approach, results are very sensitive to the weighting scheme of the loss function (Nichol and Dhariwal, 2021; Hang et al., 2023). Our interval training approach has no such weighting, since each interval is trained separately, and uniform weighting on the interval is sufficient in all cases. Thus, our algorithm does not require time-consuming tuning of the loss weighting, which is a significant advantage.
@@
Figure 5. Interval training clearly improves the visual results, especially the finer details. Associated videos are found in the supplementary materials.
4.6. Limitations and failure cases
Our method relies on a convolutional architecture which has a limited receptive field, especially in the temporal dimension. Our method may have trouble reproducing still background for long periods of time. This limitation is most visible for regions that are never seen in the video. We show an example in Figure 6 where the region inpainted changes from grass to water over a long occlusion (more than 30 frames apart) when it should stay stable.
Another limitation of our method is that each model is “single-use”, in the sense that it is adapted to a single video. This makes sense in the case of video inpainting, and greatly reduces network size and training time but it means that the model cannot be re-used, as it can be in other approaches. The optical flow is much more generalisable from one video to the next but, as we have seen, is limiting for elements such as dynamic textures. Thus, future research would involve creating a network which can be used on different videos.
Finally, our method is fully internal and thus relies on content from the video itself. Thus, it may have difficulty introducing completely new content in the video. Adding such completely new content, while maintaining a reasonable network size is a significant challenge for future research.
@@ 4.6. Limitations and failure cases
Figure 6. When generating content, our method may not guarantee long term temporal consistency, especially for textures.
5. Conclusion
We have proposed a diffusion-based method for video-inpainting. To the best of our knowledge, this is the first purely diffusion-based approach to the video inpainting problem. This is made possible thanks to our novel training scheme and a lightweight network (around 500k parameters vs several billions for most video synthesis methods using diffusion). In particular, our training method leverages only the information contained in the current video, which is sufficient to achieve high quality inpainting, as confirmed by very satisfying results both quantitatively and qualitatively. We show a dramatic improvement for dynamic textures over current methods.
A future research direction would be to handle all situations, i.e. extreme camera motion and dynamic/complex scenes, within the same method. A second direction would be to accelerate training and inference times. To do so we can draw inspiration from Latent Diffusion Models (Rombach et al., 2022), distillation techniques (Salimans and Ho, 2022), or multi-scale diffusion models (Zhou et al., 2023b). Finally a major challenge is to extend the method to situations where the inpainted region should contain an element that does not resemble anything visible in the non-occluded part of the sequence.
@@ Acknowledgments
This study has been carried out with partial financial support from the French Research Agency through the PostProdLEAP project (ANR-19-CE23-0027-01), and through the MISTIC project (ANR-19-CE40-005). Nicolas Cherel is supported by the Institut Mines Télécom, Fondation Mines-Télécom, and l’Institut Carnot TSN through the “Futur et Ruptures” grant.

### v4

@@ 3.3 Single Video (Internal) Training
Internal training vs fine-tuning. Internal training has the advantage of reducing the complexity of the training problem, but the drawback is that we are in essence overfitting on the video data. Thus, a natural question is whether a generic model can be fine-tuned for the inpainting of a specific video. This could be done using existing video inpainting models [LLQ22, ZLCC23] or video generation models [HSG22]. We carried out experiments to explore these options, and found that neither options were immediately promising. Fine-tuning a video inpainting model in fact lead to worse perceptual metrics. The large generative diffusion model generated unconvincing results in the inpainting setting. We refer the reader to the appendix for details.
3.4 Interval training
@@
To circumvent this problem, we propose to divide up the timesteps into subsets, which we call intervals. We use one lightweight network, trained on one interval at a time. The training starts from the end of the diffusion process on a first interval. Once training is done on this interval, we use the model to perform the inference on the test data for all timesteps of the interval, up to the beginning of the next interval. The model is then trained on the next time interval. This is carried out until we have reached time step $t=0$. We can, of course, save the intermediate models if we wish, but this is not really necessary for video inpainting, since the model can only be used for the current video.
@@
Table 1: Inpainting metrics on dynamic texture datasets.
@@
LPIPS ↓ SVFID ↓ PSNR ↑ SSIM ↑ Diversity ↑ LPIPS ↓ SVFID ↓ PSNR ↑ SSIM ↑
Patch [NAF14] 0.051 0.227 29.47 0.954 0.053 0.0180 0.055 36.52 0.988
ISVI [ZFL22c] 0.062 0.500 27.41 0.945 0 0.0302 0.248 34.20 0.982
E2FGVI [LLQ22] 0.046 0.288 31.58 0.961 0 0.0145 0.078 38.76 0.991
FGT [ZFL22b] 0.046 0.202 30.88 0.956 0 0.0144 0.059 37.99 0.990
ProPainter [ZLCC23] 0.046 0.261 31.77 0.960 0 0.0137 0.051 39.62 0.991
@@
This divide-and-conquer strategy has several advantages. Firstly, it allows us to achieve high quality video inpainting performance with a lightweight diffusion network, since each network specializes only one timestep interval. Secondly, we can avoid complicated weighting schemes for the loss function: in one interval, all timesteps have roughly the same weight and the weighting term can thus be set to $1$. This is not the case when one network is used and must perform well on all $T$ timesteps [HGL23, ND21].
4 Experiments
4.1 Evaluation and masks
We compare our method with several state of the art methods. We include a classical patch-based approach [NAF14] which is also internal and well adapted to the inpainting of dynamic textures, as well as more recent deep learning approaches: E2FGVI [LLQ22], ISVI [ZFL22c], FGT [ZFL22b], and ProPainter [ZLCC23]. Compared to our approach, these methods are not trained or fine-tuned on the test video, but this does not always bring better performance. Details and discussion can be found in the appendix.
We evaluate our method on complex videos featuring dynamic textures, large occlusions, and moving objects. These include, firstly, video sequences with a segmented object to be removed, from [NAF14, GTK12, LAGM19], as well as some new sequences also with segmented objects. Secondly, we use dynamic texture databases: [TBD18], featuring 59 textures, and a subset of DTDB [HW18] containing one random video from each of the 45 appearance categories. In these last two databases, we use synthetic masks, since there are no meaningful objects (people walking, etc) to be removed.
@@ 4.1 Evaluation and masks
Figure 2: In these scenes, our method satisfactorily inpaints dynamic textures preserving sharp details (top). It also better handles complex occlusions such as people crossing paths.
4.2 Training details
@@
For training, we generate masks using random strokes similarly to [KWLK19]. Training masks may overlap with the test masks, in which case no loss is actually computed in the intersection.
We use the Adam optimizer with an initial learning rate of 1e-4. We train for 200k steps for object removal, this takes 15 hours. For dynamic texture inpainting which is of lower time and spatial resolution, we train for 40k steps (3 hours). For interval training, the 1000 timesteps are divided into 20 intervals of length 50. The same training budget is allocated to each interval.
4.3 Quantitative results
We quantitatively evaluate the methods using Peak Signal-to-Noise Ratio (PSNR), structure similarity measure (SSIM), Learned Perceptual Image Patch Similarity [ZIE18] (LPIPS) and Single Video Fréchet Inception Distance (SVFID) [WLZ18, GBW20]. Finally, we also compute a diversity score following [SDM19]. As explained above, the evaluation is conducted on the two databases of dynamic textures.
The first three metrics (PSNR, SSIM and LPIPS) are computed frame-by-frame and therefore do not reflect motion discontinuities. SVFID, on the other hand, takes spatio-temporal features into account. For computing the SVFID, we follow [DMB21] and use the same model, trained on dynamic textures, for extracting the spatio-temporal features required for its computation.
Results are given in Table 1. We first observe that the LPIPS metric is significantly better with our method. In contrast, PSNR and SSIM are relatively good for all methods, the SSIM being slightly better for us and the PSNR better for the two deep-learning based methods [ZLCC23]and [LLQ22]. A visual inspection of the videos provided with this paper nevertheless clearly shows that such methods struggle to reproduce spatial details of the textures. This can be explained by the fact that the PSNR is well-known to be poorly adapted to fine details [BM18, SHC22]. The LPIPS metric, on the other hand, is more faithful to human perception.
@@ 4.3 Quantitative results
The diversity score shows better results of our approach and of the patch-based method from [NAF14] with respect to other approaches. In fact, all considered deep-learning methods always produce the same result, yielding a null diversity score.
@@
Table 2: Ablation study on the interval length during training/inference. Using no interval (length 1000) is significantly worse than our chosen interval size (50).
Interval length Dynamic textures (Tesfaldet)
LPIPS ↓ SVFID ↓ PSNR ↑ SSIM ↑
1000 0.042 0.145 29.54 0.960
100 0.042 0.121 29.78 0.957
50 0.035 0.116 30.27 0.964
10 0.039 0.111 30.12 0.958
1 0.043 0.137 30.10 0.958
@@
Table 3: Inpainting metrics when averaging 100 samples or using a single one on the Tesfaldet dataset. PSNR is improved by averaging but the perceptual metrics are severely degraded.
Method LPIPS ↓ SVFID ↓ PSNR ↑ SSIM ↑
Ours (avg) 0.056 0.231 30.97 0.962
Ours 0.035 0.116 30.27 0.964
Table 4: Inference times, for a 432x240 video on an NVidia V100.
Method Inference speed (second / frame)
Patch [NAF14] 5.13
ISVI [ZFL22c] 1.59
E2FGVI [LLQ22] 0.10
FGT [ZFL22b] 1.83
ProPainter [ZLCC23] 0.14
Infusion (ours) 5.07
4.4 Object removal experiments
@@
In this section, we consider the problem of object removal through several classical sequences used in previous works on video inpainting [NAF14, GTK12, LAGM19] and a few new ones. We consider the same methods as before. For each sequence, a mask is given corresponding to an object (often a person) to be removed. As already mentioned, there is no ground truth for these sequences, so that quantitative evaluation is not performed. Nevertheless, such situations correspond to real applications and a more realistic foreground-background articulation than with synthetic masks. All results can be found in the supplementary materials. One can observe a much better respect of backgrounds such as dynamic textures or occluded moving objects with our approach than with other deep-learning based ones. This is illustrated in Figure 2. On the first line, one sees that the unstable reflection of the building in the water is better reproduced by our methods than by FGT and ProPainter. The second and third rows display cases with occluded moving objects. In this case, results are blurry for ProPainter and blending artifacts are visible for FGT, while ours yields more satisfactory structures. We encourage the reader to view the videos in the supplementary materials. These videos support the claim that a diffusion framework with appropriate internal learning is a good choice for stochastic and complex content. Our method can however suffer from some artifacts as seen in the loulous sequence (see supplementary materials). A black artifact is present when the inpainting mask overlaps the video boundaries in the last frames, this is due to our simplistic (training) mask generation technique.
4.5 Training and inference times
Inference speeds in second / frame for all methods can be found in Table 4. We can see that our method is not the fastest, performing on par with the patch-based method of Newson et al. This is a known limitation of diffusion models with a large number of diffusion steps.
For training times, we cannot compare with the competing methods because the relevant information is not always available or is valid for a different hardware configuration. For our method on the dynamic texture datasets, training takes 3 hours on average for the Tesfaldet dataset and 6 hours on average for the DTDB dataset which has a higher resolution. All trainings are done using one NVidia V100. For object removal, we increase the number of training iterations, which results in a training time of around 15 hours. Training times could possibly be reduced by adopting latent diffusion models [RBL22].
4.6 Ablation study
@@
This section demonstrates the importance of interval training. We compare the baseline training (i.e. training on all timesteps) and interval training. We train a neural network with the same number of parameters and for the same total number of training steps for both approaches. Two visual results can be seen in Figure 3. We observe that interval training clearly improves the reproduction of fine, textured details. More results can be seen on the videos in the supplementary materials.
We also evaluate the impact of the interval length size quantitatively. Results can be found in Table 2 and indicates that there is an optimal interval length for $l=50$ which is the default parameter we use in the experiments. Note that all these experiments are done with the same total training budget. Interval training significantly improves the quantitative results, especially for the perceptual metrics (LPIPS, SVFID).
@@
We recall that in the baseline approach, results are very sensitive to the weighting scheme of the loss function [ND21, HGL23]. Our interval training approach has no such weighting, since each interval is trained separately, and uniform weighting on the interval is sufficient in all cases. Thus, our algorithm does not require time-consuming tuning of the loss weighting, which is a significant advantage.
@@
Figure 3: Interval training clearly improves the visual results, especially the finer details. Associated videos are found in the supplementary materials.
4.7 Limitations and failure cases
Table 5: Inpainting metrics on the DAVIS dataset evaluation set.
Method PSNR (dB) ↑ SSIM ↑
ISVI [ZFL22c] 30.34 0.946
E2FGVI [LLQ22] 33.71 0.970
ProPainter [ZLCC23] 34.43 0.973
Infusion (ours) 31.55 0.975
Our method relies on a convolutional architecture which has a limited receptive field, especially in the temporal dimension. Our method may have trouble reproducing still background for long periods of time. This particularly degrades the performance for static background reconstruction as shown by the comparison with other methods on the DAVIS dataset (Table 5). On this dataset, which contains mostly still background, the positioning of our method is less favorable compared to other methods. This limitation is most visible for regions that are never seen in the video. We show an example in Figure 4 where the region inpainted changes from grass to water over a long occlusion (more than 30 frames apart) when it should stay stable.
Another limitation of our method is that each model is “single-use”, in the sense that it is adapted to a single video. This makes sense in the case of video inpainting, and greatly reduces network size and training time but it means that the model cannot be re-used, as it can be in other approaches. The optical flow is more generalizable from one video to the next but, as we have seen, is limiting for elements such as dynamic textures.
Finally, our method is fully internal and thus relies on content from the video itself. Thus, it may have difficulty introducing completely new content in the video. An example of this situation can be seen in the Appendix where an object only half-occluded is not semantically inpainted. Adding such completely new content, while maintaining a reasonable network size is a significant challenge for future research.
@@ 4.7 Limitations and failure cases
Figure 4: When generating content, our method may not guarantee long term temporal consistency, especially for textures.
5 Conclusion
We have proposed a diffusion-based method for video-inpainting. This is made possible thanks to our novel training scheme and a lightweight network (around 500k parameters vs several billions for most video synthesis methods using diffusion). In particular, our training method leverages only the information contained in the current video, which is sufficient to achieve high quality inpainting, as confirmed by very satisfying results both quantitatively and qualitatively. We show a dramatic improvement for dynamic textures over current methods.
A future research direction would be to handle all situations, i.e. extreme camera motion and dynamic/complex scenes, within the same method. A second direction would be to accelerate training and inference times. To do so we can draw inspiration from Latent Diffusion Models [RBL22], distillation techniques [SH22], or multi-scale diffusion models [ZYY23]. Finally a major challenge is to extend the method to situations where the inpainted region should contain an element that does not resemble anything visible in the non-occluded part of the sequence.
@@ Acknowledgments
This study has been carried out with partial financial support from the French Research Agency through the projects PostProdLEAP (ANR-19-CE23-0027-01), MISTIC (ANR-19-CE40-005), and IDeGeN (ANR-21-CE23-0024). Nicolas Cherel is supported by the Institut Mines Télécom, Fondation Mines-Télécom, and l’Institut Carnot TSN through the “Futur et Ruptures” grant.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
FGT E2FGVI ProPainter Infusion
Figure 5: On all these challenging textures, our method better inpaints the missing region. The motions are also much better inpainted (see video in the supplementary materials).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
FGT ProPainter Infusion
Figure 6: In these scenes, we already see in static images that our method significantly outperforms other approaches which are either blurry or contain network artifacts. The result is even more clear when played as a video as optical-flow based methods are unable to reproduce the convincing dynamics of the textures.
E2FGVI Refer to caption Refer to caption Refer to caption Refer to caption
Infusion Refer to caption Refer to caption Refer to caption Refer to caption
FGT Refer to caption Refer to caption Refer to caption Refer to caption
Infusion Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Examples from DTDB. Our method is able to inpaint highly dynamic textures which is a current limitation of flow-based methods like E2FGVI, or FGT.
Refer to caption Refer to caption Refer to caption
Average of 100 samples Sample # 1 # Sample 2
Figure 8: Average of 100 diffusion samples vs independent samples. Averaging leads to a better PSNR (22.78 dB vs 20.66 dB vs 20.60 dB) but degrades the results visually. For stochastic textures, the PSNR is not a relevant metric.
Appendix A Fine-tuning on test videos
Fine-tuning is now a common practice with the availability of large foundation models [RBL22] and efficient techniques such as LoRA [HysW22]. While fine-tuning sounds more reasonable than training a model from scratch for each video, we show in this section that this does not always generate good results.
First, we show that fine-tuning does not bring additional performance for ProPainter [ZLCC23]. ProPainter indeed is trained on a separate external dataset and could benefit from a refinement step on a specific video. However we see in Table 6), that training ProPainter for an additional 200 steps (20 min) improves the PSNR but degrades the perceptual metrics such as LPIPS and SVFID. Note that ProPainter could only be fine-tuned using the L2 loss and not the discriminator loss. The discriminator cannot be used when only one partially masked video is available as training data.
Table 6: Inpainting results after fine-tuning on the Tesfaldet texture dataset. Fine-tuning ProPainter on the test video degrades the perceptual metrics (LPIPS, SVFID) but improves the PSNR. FT=fine-tuned
Method LPIPS ↓ SVFID ↓ PSNR↑ SSIM ↑
Ours 0.035 0.116 30.27 0.964
ProPainter 0.046 0.261 31.77 0.960
ProPainter FT* 0.053 0.367 32.11 0.962
Second, we tried to use one of the few video diffusion models available: Stable Video Diffusion [BDK23]. Stable Video is however not designed to do video inpainting, so several adjustments have to be made. Specifically, we adopted the approach of Lugmayr et al. in RePaint [LDR22] by removing the conditioning information and performing several time travels while reprojecting the visible regions onto the generated video. Unfortunately, this lead to poor results with unrealistic completions lacking coherence with the visible part of the video.
Table 7: U-Net Architecture Details (4 levels) - All convolutions are 3x3x3 and use ReLU.
Layer Type Output Shape (T, H, W, C) Extra
Input Concat (T, 256, 256, 23) $\tilde{x}_{t},y,t,M$
Conv1_0 Conv3D (T, 256, 256, 32) -
Conv1_1 Conv3D (T, 256, 256, 32) -
Conv1_2 Conv3D (T, 256, 256, 32) $\rightarrow$ Skip1
Pool1 MaxPool2D (T, 128, 128, 32) -
Conv2_1 Conv3D (T, 128, 128, 32) -
Conv2_2 Conv3D (T, 128, 128, 32) $\rightarrow$ Skip2
Pool2 MaxPool2D (T, 64, 64, 32) -
Conv3_1 Conv3D (T, 64, 64, 32) -
Conv3_2 Conv3D (T, 64, 64, 32) $\rightarrow$ Skip3
Pool3 MaxPool2D (T, 32, 32, 32) -
Conv4_1 Conv3D (T, 32, 32, 32) -
Conv4_2 Conv3D (T, 32, 32, 32) -
Up3 Upsample (T, 64, 64, 32) -
Concat Concat (T, 64, 64, 64) Up3, Skip3
Conv5_1 Conv3D (T, 64, 64, 32) -
Conv5_2 Conv3D (T, 64, 64, 32) -
Up2 Upsample (T, 128, 128, 32) -
Concat Concat (T, 128, 128, 64) Up2, Skip2
Conv6_1 Conv3D (T, 128, 128, 32) -
Conv6_2 Conv3D (T, 128, 128, 32) -
Up1 Upsample (T, 256, 256, 32) -
Concat Concat (T, 256, 256, 64) Up1, Skip1
Conv7_1 Conv3D (T, 256, 256, 32) -
Conv7_2 Conv3D (T, 256, 256, 32) -
Output Conv3D (T, 256, 256, 3) -
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Interval 1 Interval 10 Interval 50 Interval 100
Figure 9: Inpainting visual results with different interval sizes. Short intervals have a more noisy texture than long intervals. Best seen with zoom-in. All methods are trained with the same training budget.
Figure 10: Example of failure case. Semantic inpainting is often impossible due to our learning strategy and limited training data; we assume similarity between inpainted content and visible content. In this case, the rhino’s head is not correctly reconstructed.